OpenStudy (anonymous):

Find the value of k so that the line through the given points has slope m. (K,k+1), (3,2); m=3

OpenStudy (nikato):

WELCOME TO OPENSTUDY!!

OpenStudy (nikato):

remember \[slope/m=\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\]

OpenStudy (nikato):

plug in your values \[3=\frac{ (k+1)-2 }{ k-3 }\]

OpenStudy (nikato):

simplify the numberator \[3=\frac{ k-1 }{ k-3 }\]

OpenStudy (nikato):

multiply both sides by k-3 so 3(k-3)=k-1 distribute 3k-9=k-1 subtract k on both sides 2k-9=-1 add 9 to both sides 2k=8 divide both sides by 2 k=?